# Quasi-likelihood inference for SDE with mixed-effects observed at high frequency

### v1

@@ 4. Numerical experiments
Table 4. Simulations under Model 2: empirical mean and empirical standard error (in brackets) of $500$ the drift parameters estimates for each combination of $N$, $T$ and $n$. The true parameter values are $\mu_{1}=2$, $\mu_{2}=1$, $\omega_{1}^{2}=1$.
$(N,T)$ $n$ $\widehat{\mu_{1}}$ $\widehat{\mu_{2}}$ $\widehat{\omega_{1}^{2}}$
@@
Table 5. Simulations under Model 3: empirical mean and empirical standard error (in brackets) of $500$ the diffusion parameters estimates for each combination of $N$, $T$ and $n$. The true parameter values are $\eta=0.5$, $\lambda=1$. $\widehat{\eta_{y}}$, $\widehat{\lambda_{y}}$ stand for the estimates computed from the trajectories observations, whereas $\widehat{\lambda_{\tau}}$ denote the estimates computed from the simulated $\tau_{i}$’s.
@@
Table 6. Simulations under Model 3: empirical mean and empirical standard error (in brackets) of $500$ the drift parameters estimates for each combination of $N$, $T$ and $n$. The true parameter values are $\mu_{1}=2$, $\mu_{2}=1$, $\omega_{1}^{2}=1$, $\omega_{2}^{2}=0.5$, $\omega_{3}=0.2$.
$(N,T)$ $n$ $\widehat{\mu_{1}}$ $\widehat{\mu_{2}}$ $\widehat{\omega_{1}^{2}}$ $\widehat{\omega_{3}}$ $\widehat{\omega_{2}^{2}}$
@@ 5. Application to real data
We chose to illustrate our approach using neuronal data from the R package mixedsde. The data set consists of $N=240$ membrane potential trajectories that contain $n=2000$ points spaced by a time step of $h=0.00015$ s. This data set has been the subject of several published analyses (see e.g. [9]), in which the authors primarily employed Ornstein-Uhlenbeck processes with Gaussian random effects and fixed effects in the drift, along with a constant diffusion coefficient.
The strength of our approach lies in its ability to incorporate random effects with different distributions in the drift and to consider both random and fixed effects in the diffusion coefficient. In particular, we do not utilize a mixed-effects Ornstein-Uhlenbeck process in this analysis; instead, we apply a more general model that includes the Ornstein-Uhlenbeck process when $\eta_{1}=0$ and which incorporates random and fixed effects on both the drift and the diffusion coefficient.
$Y_{i}(t)=\tau_{i}\int_{0}^{t}(\varphi_{i1}+\varphi_{i2}Y_{i}(t))dt+\int_{0}^{t}\sqrt{\tau_{i}(1+\eta Y_{i}(t)^{2})}dW_{i}(t)$
@@
$\varphi_{i}=(\varphi_{i1},\varphi_{i2})^{\top}=\tau\sim\mathcal{N}(\mu,\Sigma)\;,\;\tau_{i}\sim\mathcal{L}(\theta_{\tau})\;,\;i=1,\ldots,N.$
Before our analysis, we scaled the signal by multiplying the membrane potential measurements by a factor of $200$. By doing so, we aimed to make visible an inter-trajectory variability that would otherwise be challenging to discern.
In the first step of our approach, we examined the empirical distribution of the estimated parameters in the diffusion coefficients of each trajectory to select the most appropriate distribution possible. We attempted to fit several distributions: exponential, Weibull, gamma, log-normal. The QQ plots led us to choose a generalized Weibull distribution, with density
$f(\tau;\gamma,\sigma,\lambda)=\gamma\left(\frac{\tau-\lambda}{\sigma}\right)^{\gamma-1}\exp\left[-\left(\frac{\tau-\lambda}{\sigma}\right)^{\gamma}\right]$
as the most fitting for our data. The parameters are estimated at the following values: $\widehat{\eta}=-0.013$, $\widehat{\lambda}=3.699$, $\widehat{\sigma}=4.592$, $\widehat{\gamma}=3.447$, $\widehat{\mu}=(-5.056,10.344)^{\top}$, $\widehat{\Sigma}=\begin{pmatrix}0.098&0.122\\ 0.122&2.322\end{pmatrix}$.
Once all model parameters were estimated, we simulated 1000 trajectories under the estimated model to graphically assess the fit quality of our model. In Figure 4, we superimposed the simulated trajectories (in gray), the 95% prediction intervals (in green), and the neuronal data (in black). The results are quite convincing, both visually and in their coherence with the parameters estimated in previous analyses, although the observation time window is not as extensive as in the simulations presented in the previous sections. However, it is difficult to claim that we have correctly identified the best model. An additional step would involve model selection, comparing the Ornstein-Uhlenbeck model with our model; however, this is beyond the scope of the present work.
Figure 4. Simulated and observed trajectories of membrane potentials. The grey lines represent 1000 simulated trajectories generated from the estimated model, illustrating the expected variability in the system. The green shaded area indicates the 95% prediction intervals for the simulated trajectories, providing a range within which future observations are expected to fall. The black lines represent the observed neuronal data, highlighting the actual recorded trajectories of membrane potentials. This comparison allows for visual assessment of the model fit and the adequacy of the estimated parameters.
@@ 6. Discussion
The stochastic hierarchical model we consider, based on a diffusion process with both fixed and random effects in the drift and the diffusion coefficient, is well suited for analyzing repeated temporal dynamics across individuals. It provides a flexible and interpretable framework to disentangle the intrinsic variability of the process—captured through the stochastic component of the diffusion—and the inter-individual variability, modeled by treating certain parameters as random variables. Such a modeling strategy is particularly relevant for data observed in discrete time but governed by an underlying continuous-time dynamic, which is common in biomedical settings such as pharmacokinetics, neuroscience, or longitudinal physiological monitoring. One of the strengths of our approach lies in the simplicity of the estimation strategy, which is easy to implement and computationally efficient. The procedure is broadly applicable, as it accommodates a wide range of distributions for the random effects in the diffusion coefficient, as long as they are supported on the positive real line. Furthermore, we establish the consistency and asymptotic normality of the resulting estimators, which ensures sound statistical inference. A particularly appealing feature of the method is its capacity to characterize the observed variability in the data through the estimated parameters, providing insight into both individual-level dynamics and group-level heterogeneity. In addition, the framework allows the inclusion of an auxiliary process as a covariate, which can further enhance the modeling of inter-individual variability and improve explanatory power in applications. However, the proposed methodology also presents some limitations. First, it relies on high-frequency observations, which may not be available in all practical scenarios. Second, the current model does not explicitly account for measurement noise, an aspect that can be critical in many experimental or clinical data settings. One possible extension to address low-frequency sampling could involve simulating additional intermediate points via Brownian bridges [3], though this remains a partial workaround and does not fully compensate for the absence of an explicit noise model. The Python codes for this work are available upon request.
Acknowledgement
This work was partially supported by JST CREST Grant Number JPMJCR2115, Japan, and by JSPS KAKENHI Grant Numbers 23K22410 and 24K21516 (HM). We are also grateful to the INRAE MIGALE bioinformatics platform (https://migale.inrae.fr) for providing computational resources.
@@ A.1. Generic notation
  • For any positive sequence $(r_{N,n})$ and random functions $\xi_{N,n}(\vartheta)$, we will write:
@@
  • We will write $\overline{\mathfrak{m}}_{ij}$ for any random variable (possibly matrix-valued) which is $\mathcal{F}_{i,t_{j}}$-measurable and satisfies that
@@
    Similarly, the notation $\overline{\mathfrak{m}}_{i}$ will be used for any $\mathcal{F}_{i,T}$-measurable random variable such that
@@
    Obviously, it is true that $\overline{\mathfrak{m}}_{i}=n^{-1}\sum_{j}\overline{\mathfrak{m}}_{ij}$ and $\overline{\mathfrak{m}}_{i}=\overline{\mathfrak{m}}_{ij}$. Further, we generically write $\overline{\mathfrak{b}}_{ij}$ for any essentially bounded $\mathcal{F}$-measurable random variable possibly depending on the index $(i,j)$. The above variables may vary at each appearance, and so are their dimensionalities.
@@
This section is devoted to the proof of Theorem 3.2. We will handle the fixed-effect and random-effect parameters in this order.
@@ A.2.1. Fixed-effect parameter
which is maximized at $\widehat{\eta}$. By the standard $M$-estimation theory, the consistency of $\widehat{\eta}$ will follow from the uniform-in-$\eta$ limit in probability of $\mathbb{Y}_{11}$ and the suitable identifiability condition.
@@
Because of the strict convexity of $s\mapsto-\log s$ ($s>0$) and Assumption 2.2, Jensen’s inequality ensures that the random variables inside the expectation are a.s. negative for $\eta\neq\eta_{0}$. This implies that $\mathbb{Y}_{11,0}(\eta)\leq 0$ with the equality holding only for $\eta=\eta_{0}$. Hence, the consistency of $\widehat{\eta}$ follows from showing that
@@
$\sup_{i}\sup_{n}E\Bigg{[}\Bigg{|}\sqrt{n}\Bigg{(}\frac{1}{n}\sum_{j}\psi_{i,j-1}-\frac{1}{T}\int_{0}^{T}\psi_{i}(t)dt\Bigg{)}\Bigg{|}^{K}\Bigg{]}<\infty.$ (A.2)
@@
We only prove (A.3), the proofs of (A.2) and (A.4) are similar. Recalling that $h=T/n$ and that $X_{1},X_{2},\dots$ are i.i.d., we can estimate as
@@
Recall $S_{i,j-1}=S_{i,j-1}(\eta_{0})$ and pick any $K\geq 2$. Applying Lemma A.1, we readily obtain for each $\eta$ and $k=0,1$,
@@
These estimates together with the Sobolev inequality (2.7) yield
@@
$Q_{i}(\eta):=\frac{1}{n}\sum_{j}S_{i,j-1}^{-1}(\eta)\,y_{ij}^{2}.$ (A.8)
We will identify the uniform-in-$\eta$ limit in probability of the first term on the right-hand side of in (A.7). Let us write
$y_{ij}=\sqrt{\tau_{i}}\,c_{i,j-1}(\eta_{0})z_{ij}+\sqrt{h}\,\mu_{ij},$
where $z_{ij}:=h^{-1/2}\Delta_{j}w_{i}\sim\text{i.i.d.}\leavevmode\nobreak\ N(0,1)$ (with respect to both $i$ and $j$) and where
@@
$y_{ij}^{2}$ $=\tau_{i}\,S_{i,j-1}z_{ij}^{2}+\sqrt{h}\,\overline{\mathfrak{m}}_{ij}$
@@
The finer representation (A.9) is not used here, but will do matter later in the proof of the asymptotic normality.
@@
the last one does not really define $\delta_{3,i}(\eta)$ since $\overline{\mathfrak{m}}_{ij}$ is a generic notation, but that form will be enough. By Assumption 2.1, the random function $F_{i}(\eta)$ is essentially bounded and also bounded away from 0: there exists a constant $K>0$ for which
@@
Pick any $M>2\vee p_{\eta}$. By a similar argument to (A.6) in the proof of Lemma A.1 together with the Sobolev inequality, we obtain $\sup_{i,j}E[\sup_{\eta}|\delta_{1,i}(\eta)|^{M}]<\infty$. Similarly, since $n^{-1}h^{-1/2}=T^{-1/2}n^{-1/2}$, the Sobolev and Burkholder inequalities lead to $\sup_{i,j}E[\sup_{\eta}|\delta_{2,i}(\eta)|^{M}]<\infty$. The bound $\sup_{i,j}E[\sup_{\eta}|\delta_{3,i}(\eta)|^{M}]<\infty$ is trivial. It follows that
@@
Pick any sequence $(\kappa_{N})\subset(0,1/2]$ such that $\kappa_{N}\to 0$ and that $\sqrt{n}\,\kappa_{N}\gtrsim N^{\mathsf{a}_{11}}$ for some constant $\mathsf{a}_{11}>0$ (possible because of (2.11)). We will show
$D(\eta)=o_{p}^{\ast}(\kappa_{N}^{\mathsf{b}})$ (A.15)
for any $\mathsf{b}\in(0,1)$, which concludes the following estimate slightly stronger than (A.1)111 About the rate of convergence of (A.1): We can take $\kappa_{N}\sim N^{\mathsf{a}^{\prime\prime}}n^{-1/2}\lesssim n^{\mathsf{a}_{11}\mathsf{a}^{\prime\prime}-1/2}$ with $\mathsf{a}_{11}\mathsf{a}^{\prime\prime}<1/2$. Then, for any positive $\mathsf{c}\leq(1+\mathsf{a}^{\prime\prime})^{-1}\min\{\mathsf{a}^{\prime}/2,(1/2-\mathsf{a}_{11}\mathsf{a}^{\prime})\mathsf{b}\}$, we have $(nN)^{\mathsf{c}}(\mathbb{Y}_{11}(\eta)-\mathbb{Y}_{11,0}(\eta))=O_{p}^{\ast}(1).$ (A.16) This observation will help to study BIC-type statistics; see [13, Theorem A.1] and [14, Section 4]. :
@@
For any $M>\mathsf{a}_{11}^{-1}$, recalling (A.13) and applying the Markov and Cauchy-Schwarz inequalities we have for $M>0$,
@@
Now we turn to the proof of the asymptotic normality of $\widehat{\eta}$.
We begin with the asymptotic behavior of the quasi-score function
@@
$\Delta_{11}=\frac{1}{2\sqrt{N}}\sum_{i=1}^{N}\left(\tau_{i}^{-1}Q_{i}\right)^{-1}\frac{1}{\sqrt{n}}\sum_{j}\left(\frac{y_{ij}^{2}}{\tau_{i}S_{i,j-1}}-\tau_{i}^{-1}Q_{i}\right)\,g_{i,j-1}.$ (A.19)
@@
$\frac{y_{ij}^{2}}{\tau_{i}S_{i,j-1}}-\tau_{i}^{-1}Q_{i}$ $=\zeta_{ij}+\sqrt{h}\,\overline{\mathfrak{m}}_{ij},$ (A.21)
@@
by taking $K>0$ sufficiently large. In particular, this ensures that
@@
$P[|\Delta_{11}^{r,2}|>\epsilon]$ $\leq P\left[\max_{i\leq N}\left|\sqrt{h}\,\frac{\delta_{i}}{\tau_{i}}\right|>\frac{1}{2}\right]+P\left[|\Delta_{11}^{r,2}|>\epsilon,\leavevmode\nobreak\ \max_{i\leq N}\left|\sqrt{h}\,\frac{\delta_{i}}{\tau_{i}}\right|\leq\frac{1}{2}\right]$
@@
It follows that Hence $\Delta_{11}^{r,2}=o_{p}(1)$.
Turning to $\Delta_{11}^{r,3}$, we need to look at its leading terms in a more detailed manner: we trace back the manipulations (A.9), (A.2.1), and (A.21). By straightforward yet a bit messy computations, it can be seen that
@@
Below, to conclude $\Delta_{11}^{r,3}=o_{p}(1)$, we will show that $\overline{\Delta}_{11}^{r,3,1}$ and $\overline{\Delta}_{11}^{r,3,2}$ are both $o_{p}(1)$.

### v2

@@ 4. Numerical experiments
Table 4. Simulations under Model 2: empirical mean and empirical standard error (in brackets) of $500$ drift parameters estimates for each combination of $N$, $T$ and $n$. The true parameter values are $\varphi_{f}^{(1)}=2$, $\varphi_{f}^{(2)}=1$, $\sigma_{r}^{2}=1$.
$(N,T)$ $n$ $\widehat{\varphi_{f}^{(1)}}$ $\widehat{\varphi_{f}^{(2)}}$ $\widehat{\sigma_{r}^{2}}$
@@
Table 5. Simulations under Model 3: empirical mean and empirical standard error (in brackets) of $500$ diffusion parameters estimates for each combination of $N$, $T$ and $n$. The true parameter values are $\eta=0.5$, $\lambda=1$. $\widehat{\eta_{y}}$, $\widehat{\lambda_{y}}$ stand for the estimates computed from the trajectories observations, whereas $\widehat{\lambda_{\tau}}$ denotes the estimates computed from the simulated $\tau_{i}$’s.
@@
Table 6. Simulations under Model 3: empirical mean and empirical standard error (in brackets) of $500$ drift parameters estimates for each combination of $N$, $T$ and $n$. The true parameter values are $\varphi_{f}^{(1)}=2$, $\varphi_{f}^{(2)}=1$, $\sigma_{r,1}^{2}=1$, $\sigma_{r,2}^{2}=0.5$, $\sigma_{r,1,2}=-0.2$.
$(N,T)$ $n$ $\widehat{\varphi_{f}^{(1)}}$ $\widehat{\varphi_{f}^{(2)}}$ $\widehat{\sigma_{r,1}^{2}}$ $\widehat{\sigma_{r,1,2}}$ $\widehat{\sigma_{r,2}^{2}}$
@@ 5. Application to real data
We chose to illustrate our approach using neuronal data from the R package mixedsde. The data set consists of $N=240$ membrane potential trajectories that contain $n=2000$ points spaced by a time step of $h=0.00015$ s. This data set has been the subject of several published analyses (see e.g. [dion2019mixedsde]), in which the authors primarily employed Ornstein-Uhlenbeck processes with Gaussian random effects and fixed effects in the drift, along with a constant diffusion coefficient.
Because the theoretical guarantees of our estimation procedure have so far been proven only for models with bounded coefficients, we began our analysis of the neural data by focusing on two models that meet these assumptions, leaving out the Ornstein–Uhlenbeck model, which does not comply with them:
  • Neural data model 1 (N1):
    $Y_{i}(t)$ $=Y_{i}(0)+\tau_{i}\int_{0}^{t}\left(\frac{\varphi_{i1}Y_{i}(s)}{\sqrt{1+Y_{i}(s)^{2}}}+\frac{\varphi_{i2}}{\sqrt{1+Y_{i}(s)^{2}}}\right)ds$
    ${}\qquad+\sqrt{\tau_{i}}\int_{0}^{t}\exp(\eta\textrm{atan}(Y_{i}(s)))dw_{i}(s)$ (5.1)
  • Neural data model 2 (N2):
    $Y_{i}(t)$ $=Y_{i}(0)+\tau_{i}\int_{0}^{t}\left(\frac{\varphi_{i1}Y_{i}(s)}{\sqrt{1+Y_{i}(s)^{2}}}+\frac{\varphi_{i2}}{\sqrt{1+Y_{i}(s)^{2}}}\right)ds$
    ${}\qquad+\sqrt{\tau_{i}}\int_{0}^{t}\exp(\eta s/2)dw_{i}(s)$ (5.2)
with, for both (N1) and (N2)
$\varphi_{i}=(\varphi_{i1},\varphi_{i2})^{\top}\sim_{i.i.d.}N_{2}(\varphi_{f},\Sigma)\;,\;\tau_{i}\sim\mathcal{L}(\theta_{\tau})\;,\;i=1,\ldots,N.$
Before our analysis, we scaled the signal by multiplying the membrane potential measurements by a factor of $200$. This scaling was performed to make visible an inter-trajectory variability that would otherwise be challenging to discern.
In the first step of our approach, we examined the empirical distribution of the estimated parameters in the diffusion coefficients of each trajectory to select the most appropriate underlying distribution. We attempted to fit several distributions: exponential, (generalized) Weibull, gamma, and log-normal. The QQ plots indicated that a generalized Weibull distribution provided the best fit. Its density is given by
$f(\tau;\gamma,\alpha,\lambda)=\gamma\left(\frac{\tau-\lambda}{\alpha}\right)^{\gamma-1}\exp\left[-\left(\frac{\tau-\lambda}{\alpha}\right)^{\gamma}\right].$
Numerical instabilities due to a non-invertible matrix $\Sigma$ in the second step motivated us to specify the first parameter $\varphi_{i1}$ as a fixed effect and the second $\varphi_{i2}$ as a random effect with variance $\sigma_{r}^{2}$. The strength of our approach lies in its ability to incorporate random effects following different distributions in the diffusion coefficient, while allowing both random and fixed effects in the drift term. The estimated parameter values are as follows:
  • (N1): $\widehat{\eta}=-0.075$, $\widehat{\lambda}=4.123$, $\widehat{\alpha}=5.093$, $\widehat{\gamma}=3.437$, $\widehat{\varphi_{f}}=(-5.017,10.260)^{\top}$, $\widehat{\sigma_{r}^{2}}=2.444$.
  • (N2): $\widehat{\eta}=-0.273$, $\widehat{\lambda}=3.683$, $\widehat{\alpha}=4.546$, $\widehat{\gamma}=3.463$, $\widehat{\varphi_{f}}=(-5.595,11.461)^{\top}$, $\widehat{\sigma_{r}^{2}}=3.156$.
Once all model parameters had been estimated, we simulated 1000 trajectories under the estimated models to graphically assess the fit quality of models (N1) and (N2). The corresponding results are shown in Figures 4(a) and 4(b) for models (N1) and (N2), respectively. In each figure, the simulated trajectories (in gray) are displayed together with the 95% prediction intervals (in green) and the neuronal data (in black). The results appear visually convincing, despite the observation time frame ($T$) being shorter than in the simulations presented in the previous sections. However, a visual comparison alone does not allow us to reliably identify which of these two models offers a superior fit to the data.
(a) Model (N1).
(b) Model (N2).
Figure 4. Models (N1) and (N2). Simulated and observed trajectories of membrane potentials. The grey lines represent 1000 simulated trajectories generated from the estimated model, illustrating the expected variability in the system. The green shaded area indicates the 95% prediction intervals for the simulated trajectories, providing a range within which future observations are expected to fall. The black lines represent the observed neuronal data, highlighting the actual recorded trajectories of membrane potentials. This comparison allows for visual assessment of the model fit and the adequacy of the estimated parameters.
Note that we also applied a more general model, consistent with those commonly used in the literature, which includes the Ornstein–Uhlenbeck process as a special case when $\eta=0$. This model, which incorporates both random and fixed effects in the drift and diffusion coefficients, does not satisfy the assumptions required to establish theoretical properties of the estimators:
  • Neural data model 3 (N3):
    $Y_{i}(t)=Y_{i}(0)+\tau_{i}\int_{0}^{t}(\varphi_{i1}+\varphi_{i2}Y_{i}(s))ds+\int_{0}^{t}\sqrt{\tau_{i}(1+\eta Y_{i}(s)^{2})}dw_{i}(s)$
@@
    $\varphi_{i}=(\varphi_{i1},\varphi_{i2})^{\top}\sim N_{2}(\varphi_{f},\Sigma)\;,\;\tau_{i}\sim\mathcal{L}(\theta_{\tau})\;,\;i=1,\ldots,N.$
Again, a generalized Weibull distribution best fitted the data. The estimated parameter values for model (N3) are $\widehat{\eta}=-0.013$, $\widehat{\lambda}=3.699$, $\widehat{\alpha}=4.592$, $\widehat{\gamma}=3.447$, $\widehat{\varphi_{f}}=(-4.838,9.770)^{\top}$, $\widehat{\Sigma}=\begin{pmatrix}0.039&0.028\\ 0.028&2.148\end{pmatrix}$.
Visual predictive checks are shown in Figure 5. Although model (N3) lacks formal theoretical guarantees, the results suggest that the estimation procedure performs well and fits the data as well as models (N1) and (N2). A natural next step would be to perform model selection, comparing models (N1), (N2), and (N3); however, this is beyond the scope of the present work.
Figure 5. Model (N3). Simulated and observed trajectories of membrane potentials. The grey lines represent 1000 simulated trajectories generated from the estimated model, illustrating the expected variability in the system. The green shaded area indicates the 95% prediction intervals for the simulated trajectories, providing a range within which future observations are expected to fall. The black lines represent the observed neuronal data, highlighting the actual recorded trajectories of membrane potentials. This comparison allows for visual assessment of the model fit and the adequacy of the estimated parameters.
@@ 6. Discussion
The stochastic hierarchical model we consider, based on a diffusion process with both fixed and random effects in the drift and the diffusion coefficient, is well-suited for analyzing repeated temporal dynamics across individuals. It provides a flexible and interpretable framework to disentangle the intrinsic variability of the process—captured through the stochastic component of the diffusion—and the inter-individual variability, modeled by treating certain parameters as random variables. Such a modeling strategy is particularly relevant for data observed in discrete time but governed by an underlying continuous-time dynamic, which is common in biomedical settings such as pharmacokinetics, neuroscience, or longitudinal physiological monitoring. One of the strengths of our approach lies in the simplicity of the estimation strategy, which is easy to implement and computationally efficient. The procedure is broadly applicable, as it accommodates a wide range of distributions for the random effects in the diffusion coefficient, as long as they are supported on the positive real line. Furthermore, we establish the consistency and asymptotic normality of the resulting estimators, which ensures sound statistical inference. A particularly appealing feature of the method is its capacity to characterize the observed variability in the data through the estimated parameters, providing insight into both individual-level dynamics and group-level heterogeneity. It is worth noting that our framework also allows for extension by incorporating an auxiliary process as a covariate into the coefficients, which can significantly refine the modeling of inter-individual variability and improve explanatory power in applications; we have conducted simulation experiments for such a case in Model 1 in Section 4 and Model 2 in Section 5.
However, the proposed methodology also presents some limitations. First, it relies on high-frequency observations, which may not be available in all practical scenarios. Second, the current model does not explicitly account for measurement noise, an aspect that can be critical in many experimental or clinical data settings. One possible extension to address low-frequency sampling could involve simulating additional intermediate points via Brownian bridges [bladt2016simulation], though this remains a partial workaround and does not fully compensate for the absence of an explicit noise model. The Python codes for this work are hosted on GitHub (https://github.com/madelattre/Paper----QL-MSDE-HF----Python-codes).
@@ A.1. Generic notation
  • We write
    $\sup_{i}=\sup_{N\geq 1}\max_{i\leq N},\qquad\sup_{j}=\sup_{n\geq 1}\max_{j\leq n},\qquad\sup_{i,j}=\sup_{n,N\geq 1}\max_{i\leq N,\,j\leq n},$ (A.1)
    where the supremum $\sup_{n,N\geq 1}$ is taken under (2.11).
  • For any positive sequence $(r_{N,n})$ and random functions $\xi_{N,n}(\vartheta)$, we write:
@@
  • We write $\overline{\mathfrak{m}}_{ij}$ for any random variable (possibly matrix-valued) which is $\mathcal{F}_{i,t_{j}}$-measurable and satisfies that
@@
    Similarly, the notation $\overline{\mathfrak{m}}_{i}$ is used for any $\mathcal{F}_{i,T}$-measurable random variable such that
@@
    Obviously, it is true that $\overline{\mathfrak{m}}_{i}=n^{-1}\sum_{j}\overline{\mathfrak{m}}_{ij}$ and $\overline{\mathfrak{m}}_{i}=\overline{\mathfrak{m}}_{ij}$. Further, we generically write $\overline{\mathfrak{b}}_{ij}$ for any essentially bounded $\mathcal{F}$-measurable random variable possibly depending on the index $(i,j)$:
    $\max_{i,j}|\overline{\mathfrak{b}}_{ij}|<\infty,\qquad\text{a.s.}$ (A.2)
    The above variables may vary at each appearance, and so are their dimensionalities.
@@
This section is devoted to the proof of Theorem 3.2. We will handle the fixed-effect and random-effect parameters step by step in this order, which is possible due to the mixed-rates structure of the first-stage profile quasi-likelihood $\mathbb{H}_{1}(\vartheta_{1})$.
@@ A.2.1. Fixed-effect parameter
which is maximized at $\widehat{\eta}$. By the standard $M$-estimation theory, the consistency of $\widehat{\eta}$ will follow from the uniform-in-$\eta$ convergence in probability of $\mathbb{Y}_{11}$ and the suitable identifiability condition.
@@
Because of the strict convexity of $s\mapsto-\log s$ ($s>0$) and Assumption 2.2, Jensen’s inequality ensures that the random variables inside the last expectation are a.s. negative for $\eta\neq\eta_{0}$. This implies that $\mathbb{Y}_{11,0}(\eta)\leq 0$ with the equality holding only for $\eta=\eta_{0}$. Hence, the consistency of $\widehat{\eta}$ follows from showing that
@@
$\sup_{i}\sup_{n}E\Bigg[\Bigg|\sqrt{n}\Bigg(\frac{1}{n}\sum_{j}\psi_{i,j-1}-\frac{1}{T}\int_{0}^{T}\psi_{i}(t)dt\Bigg)\Bigg|^{K}\Bigg]<\infty,$ (A.4)
@@
We only prove (A.5), the proofs of (A.4) and (A.6) are similar. Recalling that $h=T/n$ and that $Y_{1},Y_{2},\dots$ are i.i.d., we can estimate as
@@
We will use the following form of the Sobolev inequality (see, for example, [Ada73, Theorem 1.4.2]): given a bounded convex domain $\mathsf{D}\subset\mathbb{R}^{s}$ for some $s\geq 1$ and for any $\mathcal{C}^{1}(\mathsf{D})$-function $\psi$ and a constant $\alpha>s$, we have
$\sup_{y\in\mathsf{D}}|\psi(y)|^{\alpha}\leq C\int_{\mathsf{D}}\left(|\psi(y)|^{\alpha}+|\partial_{y}\psi(y)|^{\alpha}\right)dy$ (A.9)
for some universal constant $C=C(\mathsf{D},\alpha)>0$.
Recall the shorthand $S_{i,j-1}=S_{i,j-1}(\eta_{0})$ and pick any $K\geq 2$. Applying Lemma A.1, we readily obtain for each $\eta$ and $k=0,1$,
@@
These estimates together with (A.9) yield
@@
$Q_{i}(\eta):=\frac{1}{n}\sum_{j}S_{i,j-1}^{-1}(\eta)\,\overline{y}_{ij}^{2}.$ (A.11)
We will identify the uniform-in-$\eta$ limit in probability of the first term on the right-hand side of (A.10). Let us write
$\overline{y}_{ij}=\sqrt{\tau_{i}}\,c_{i,j-1}(\eta_{0})z_{ij}+\sqrt{h}\,\mu_{ij},$
where $z_{ij}:=h^{-1/2}\Delta_{j}w_{i}\sim\text{i.i.d.}\penalty 10000\ N(0,1)$ (with respect to both $i$ and $j$) and where
@@
$\overline{y}_{ij}^{2}$ $=\tau_{i}\,S_{i,j-1}z_{ij}^{2}+\sqrt{h}\,\overline{\mathfrak{m}}_{ij}$
@@
The finer representation (A.12) is not used here, but will matter later in the proof of the asymptotic normality.
@@
the last one does not really define $\delta_{3,i}(\eta)$ since $\overline{\mathfrak{m}}_{ij}$ is a generic notation, but that form will be enough. By Assumption 2.1, the random function $F_{i}(\eta)$ is essentially bounded and also there exists a constant $K>0$ for which
@@
Pick any $M>2\vee p_{\eta}$. By a similar argument to (A.8) in the proof of Lemma A.1 together with the Sobolev inequality (A.9), we obtain $\sup_{i,j}E[\sup_{\eta}|\delta_{1,i}(\eta)|^{M}]<\infty$. Similarly, since $n^{-1}h^{-1/2}=T^{-1/2}n^{-1/2}$, the Sobolev and Burkholder inequalities lead to $\sup_{i,j}E[\sup_{\eta}|\delta_{2,i}(\eta)|^{M}]<\infty$. The bound $\sup_{i,j}E[\sup_{\eta}|\delta_{3,i}(\eta)|^{M}]<\infty$ is trivial. It follows that
@@
Pick any sequence $(\kappa_{N})\subset(0,1/2]$ such that $\kappa_{N}\to 0$ and that $\sqrt{n}\,\kappa_{N}\gtrsim N^{\mathsf{a}_{11}}$ for some constant $\mathsf{a}_{11}>0$ (possible because of (2.11)). We will show that
$\forall\mathsf{b}\in(0,1)\quad D(\eta)=o_{p}^{\ast}(\kappa_{N}^{\mathsf{b}}),$ (A.18)
which concludes the following estimate slightly stronger than (A.3):
@@
For any $M>\mathsf{a}_{11}^{-1}$, recalling (A.16) and applying the Markov and Cauchy-Schwarz inequalities, we have for $M>0$,
@@
Now we turn to the proof of the asymptotic normality of $\widehat{\eta}$. First, we study the asymptotic behavior of the quasi-score function
@@
$\Delta_{11}=\frac{1}{2\sqrt{N}}\sum_{i=1}^{N}\left(\tau_{i}^{-1}Q_{i}\right)^{-1}\frac{1}{\sqrt{n}}\sum_{j}\left(\frac{\overline{y}_{ij}^{2}}{\tau_{i}S_{i,j-1}}-\tau_{i}^{-1}Q_{i}\right)\,g_{i,j-1}.$ (A.21)
@@
$\frac{\overline{y}_{ij}^{2}}{\tau_{i}S_{i,j-1}}-\tau_{i}^{-1}Q_{i}$ $=\zeta_{ij}+\sqrt{h}\,\overline{\mathfrak{m}}_{ij},$ (A.23)
@@
by taking $K>0$ sufficiently large. In particular, (A.30) ensures that
@@
$P[|\Delta_{11}^{r,2}|>\epsilon]$ $\leq P\left[\max_{i\leq N}\left|\sqrt{h}\,\frac{\delta_{i}}{\tau_{i}}\right|>\frac{1}{2}\right]+P\left[|\Delta_{11}^{r,2}|>\epsilon,\penalty 10000\ \max_{i\leq N}\left|\sqrt{h}\,\frac{\delta_{i}}{\tau_{i}}\right|\leq\frac{1}{2}\right]$
@@
Hence $\Delta_{11}^{r,2}=o_{p}(1)$.
Turning to $\Delta_{11}^{r,3}$, we need to look at its leading terms in a more detailed manner. We trace back the manipulations (A.12), (A.2.1), and (A.23). By straightforward yet a bit messy computations, it can be seen that
@@
To conclude $\Delta_{11}^{r,3}=o_{p}(1)$, we will show that $\overline{\Delta}_{11}^{r,3,1}$ and $\overline{\Delta}_{11}^{r,3,2}$ are both $o_{p}(1)$.